# CoMix: A Comprehensive Benchmark for Multi-Task Comic Understanding

### v1

@@ 1 Introduction
In recent works [19, 27], authors approached comics research to more complex tasks moving from detection to diarization, where low-level tasks - such as detecting objects, defining reading order and speaker identification and character re-identification - are used as a medium to generate an ordered transcription of who said what on single page. Despite the innovative application of this complex bottom-up approach to comics understanding tasks, there is no benchmark on character naming or diarization, nor metrics to assess their correctness.
@@
CoMix contains 3.8k images, gathered from almost 100 books, densely annotated with 130K objects across the four classes considered, 30k text-characters links, with 33k characters clusters with 16k total names identified. An overview of single-page annotations is provided in Figure 2 while statistics about datasets and annotations are provided in Table 2. We open-source the images and annotations of the validation splits. An evaluation server and images from the held-out test split are made available. Since currently there is no model that can tackle all the evaluation tasks in our benchmark, we provide baseline results for per-task models: object detection, speaker identification, character re-identification, reading order, character naming, and dialog generation.
@@ 3 Books in the CoMix
The CoMix dataset has been meticulously curated to showcase a broad spectrum of comic book styles, drawing samples from Japanese manga (PopManga), American comics (DCM and COMCIS), and French Bandes Dessinées (eBDtheque). Although multilingualism is not a primary objective, some pages from eBDtheque include French and Japanese, reflecting incidental multilingual aspects.
@@
Using this information, we selected comic books based on the distribution of character appearances across the books following the principle of (i) most characters possible, from the minimum number of books, and (ii) all possible instances of these characters should be in the 100 books selected. In constructing CoMix, our primary goal is to underscore books featuring characters that recurrently appear across various publications, indicating significant narrative roles. To achieve this, the selection method computes the ratio of shared-to-unique character appearances for each book, thereby ranking and choosing the top books based on this ratio. Specifically, this ratio assesses the frequency of characters’ appearances within the same book against their appearances in other books. To further refine our selection, the algorithm was designed to maximize the diversity of characters across the selected books, prioritizing the retention of books that uniquely feature these characters. This ensures a broad representation of characters, with most being exclusive to the selected subset and not featured in the broader set of $22k$ comics. The ratio is then exponentiated (in our implementation, squared) to enhance distinctions between books with high and low levels of character sharing. After calculating these ratios for all books, they are ranked in descending order based on their scores, leading to the selection of the top 100 books. This metric, referred to as the “Pow Selection Approach”, is detailed in the Appendix A. This approach ensures that the selected books reflect broader narrative arcs and character interconnectivity, thereby enriching the CoMix dataset’s utility for diverse comic book analysis.
Splits: The CoMix contains 85 books from different sources, each of 44 pages on average, with a total of 3.8k images. This corresponds only to the held-out test split available through the evaluation server. Moreover, a validation split (10 books, 500 images), on which we tested and reported the results in this paper is provided.
@@ 4 Annotation in the CoMix benchmark
Table 2: Annotations in the CoMix benchmark. The reported numbers correspond to both test (80 books) and validation (20 books) splits. Each detected object is represented by the class and bounding box. Text elements also possess GT transcriptions. Character elements, whenever relevant, are associated with the name. Not all texts are associated with a character, but all the texts have Reading order GT.
@@
Dialog Generation. As a way to connect high-level and low-level single-page comics understanding capabilities, we define dialogs as an ordered list of tuples with the speaker name (character names or “narrator”). The texts are sorted based on the GT reading order. Annotators are instructed to maintain punctuations and text format (lower-case and upper-case) but discard other text properties (bold, italics, canceled, or underlined text).
@@ 5 Baselines results
Computational tasks: We defined six computational tasks based on the annotations available in the CoMix dataset. We summarise in Table 3 the task definitions (outputs, metrics) and the performance of top-performing baselines. It can be observed that the CoMix combines lower-level dense prediction tasks like object, speaker identification, and character re-identification, whose outputs are box and group of indexes, with higher-level tasks like dialog generation. More details about the task definitions are included in the appendix.
Baselines: Ideally, a single model should be able to perform all the tasks in the CoMix benchmark. Since such a model is not available in the literature, we include results obtained with per-task baselines on the validation split for all the six tasks in the CoMix; see Table 3 for a summary of top-performing baselines and their average performance, and the appendix for more details. When selecting and running these baselines, we favored the same approaches used by Sachdeva and Zisserman [27], which we carefully detail in the following section. However, for character naming, such models do not exist in the literature, so we evaluated namings together with dialogs, instead of with detection and clustering.
@@
Object detection: For object detection, we selected two common metrics: mean Average Precision at IoU of $0.5$ ($mAP@0.5$) and Recall at 100 objects ($R@100$). We benchmarked a variety of models including convolutional and transformer-based architectures, with a focus on their performance in both fine-tuned and zero-shot settings on comic data. Among these, GroundingDino [20], designed for open-set object detection using natural language object classes, was a key zero-shot model utilized to detect four classes: panels, characters, text, and faces. Additionally, conventional models such as Faster R-CNN, SSD, and YOLO were trained on varied comic styles to assess the impact of different training data distributions. For characters and faces detection we employed also a YOLOX-based available model named DASS [31]. Notably, MAGI [27], a transformer-based model, was included for its impressive capability demonstrated in Manga-style comics. Details about the fine-tuned training procedure and GroundingDino prompts for zero-shot detection are provided in the appendix B, as well as more detailed benchmark results.
@@
Character Naming & Dialog Generation. In addressing the challenges of character naming and dialog generation, we acknowledge that there is a lack of specific metrics. In previous works, the dialog is generated with a combination of algorithmic panel ordering, character progressive naming (“character 1”, “character 2”), and fine-tuned OCR for textbox transcription. This engineering and multi-step approach is a good showcase but is not evaluated with some metric against any ground truth. We introduce a smoothed case-sensitive edit distance metric called “Hybrid Dialog Score” ($HDS$) to evaluate the accuracy of model outputs against ground-truth dialog annotations. This metric assesses both the precision of transcribed dialogues and the accuracy of character identification in a unified framework. The metric operates in three steps: (i) we match gt and predictions texts using Hungarian matching with edit distance metric; (ii) for every match, we calculate the textbox edit distance (normalized on ground truth text length); (iii) for every match, we calculate the character name ANLS. An in-depth overview of the metric is provided in the supplementary materials, together with pseudocode. For our benchmarks, we utilized MAGI as described in [27]. As MAGi is not able to detect names, we employ GPT-4 [25], leveraging its multimodal capabilities to interpret comic pages and generate structured outputs that include both dialogues and character names. GPT-4 was tasked with identifying and naming characters, where known, or assigning incremental identifiers otherwise, and transcribing dialogues exactly as they appear in comic format, respecting the sequence and case sensitivity. This approach not only captures the complexity of comic narratives but also enhances the evaluation of dialog transcription fidelity and character consistency across various comic styles. The result is reported in Table 3, and an overview of the GPT-4 prompt to obtain structured predictions is given in supplementary materials.
@@ 6 Conclusion
In this paper, we introduced CoMix, a comprehensive benchmark designed to evaluate multi-task and multi-modal capabilities in comic analysis. Addressing the limitations of existing datasets, CoMix incorporates diverse comic styles, including American and manga, and offers extensive annotations for a wide range of tasks. This benchmark sets a new standard for evaluating comic analysis models by providing a robust and varied test bed.
Our contributions include identifying the lack of comprehensive benchmarks and metrics, introducing the Hybrid Dialog Score for evaluating character naming and dialog generation, and presenting CoMix as a diverse dataset with annotations across multiple tasks and styles. Baseline results reveal significant performance gaps between current models and human performance, highlighting the challenges and opportunities for future advancements.
The public release of the CoMix validation split and an evaluation server for the held-out test split encourages open research and benchmarking. Future work should focus on expanding the dataset to include more diverse languages. By providing a comprehensive and challenging benchmark, we aim to inspire the research community to advance the understanding and analysis of comics, bridging the gap between human and machine comprehension in this culturally rich medium.
Supplementary Materials
@@
This supplementary document complements the main paper by providing additional information and examples that could not be included within the page constraints of the original manuscript. The structure of this document is as follows: in Section A we describe the book selection while in Section B details about the models and their usage. In Section C we presents an overview of the annotations from existing datasets and compares them with those in the CoMix benchmark. In Section D we detail the selection process of American Comics books included in our study. Extensive results are provided in Section E, including those reported in the main manuscript. Finally, copyright information is outlined in Section F. This supplementary material is intended to enhance the understanding and transparency of the research presented in the main paper.
Appendix A Book Selection
1. Calculation of Shared-to-Unique Character Ratio: For each book, a ratio is calculated based on the frequency of characters appearing within and in other books. This ratio is then exponentiated (typically squared) to emphasize differences between books regarding character sharing.
Algorithm 1 Book Selection Based on Character Sharing
1:procedure PowSelectionApproach( $book\_to\_characters$, pow)
2:     $book\_shared\_to\_unique\_ratio$leftarrow empty dictionary
3:     for $book\_id$, $characters$ in $book\_to\_characters$ do
4:         $shared\_count\leftarrow 0$
5:         $unique\_character\_count\leftarrow\text{len( }characters$ )
6:         for $other\_book\_chars$ in $book\_to\_characters.values$ do
7:              if $other\_book\_chars$neq $characters$ then
8:                  for $char\_id$ in $characters$ do
9:                       if $char\_id$ in $other\_book\_chars$ then
10:                           $shared\_count\leftarrow shared\_count+1$
11:                       end if
12:                  end for
13:              end if
14:         end for
16:         $book\_shared\_to\_unique\_ratio[book\_id]$leftarrow $ratio^{{\color[rgb]{0.4,0.6,0.8}pow}}$
17:     end for
18:     return $book\_shared\_to\_unique\_ratio$
19:end procedure
This algorithm ensures that the selected books reflect a broader narrative context, as it prioritizes those with characters that bridge multiple storylines. Such an approach is suitable for analyzing complex datasets where character interrelations are significant.
Appendix B Detailed Model Descriptions and Settings
@@
B.1 Hybrid Dialog Score
A detail of the new proposed metric for dialogs score is provided in the following algorithm.
Algorithm 2 Hybrid Dialog Score
3:     $tot\_ed$, $char\_name\_score$ $\leftarrow 0,0,0$
4:     for each $(mo,gt)$ in $matches$ do
6:         $tot\_ed$leftarrow $tot\_ed$ $+$ $edit\_dist$ /// len($gt.text$)
7:         $anls\_score\leftarrow\text{calculate ANLS}(mo.name,gt.name)$
8:         $char\_name\_score$leftarrow $char\_name\_score$ $+$ $anls\_score$
9:     end for
10:     $tot\_ed$ $\leftarrow 1-$ $tot\_ed$
12:     return $tot\_ed$, $char\_name\_score$
13:end procedure
Appendix C Annotations overview
@@
Additionally, annotations for speaker identification, character re-identification, and reading order may be less visible in the images provided. These are represented as green and red lines connecting the text to the character boxes (speaker identification), colored points identifying each character bounding box (character re-identification), and purple splines connecting the text boxes in reading order orientation (reading order). For a clearer view of these annotations, please visit the CVAT website777only available in the original submitted paper with the following credentials [^8] USER: ********** - PSW: ***********.
@@ C.2 Qualitative details
Appendix D Data selection
The CoMix benchmark sees the presence of selected American golden-age comics from DCM  [^9] Digital Comics Museum at https://digitalcomicmuseum.com. These documents have not been selected following the “most downloaded” principle, as instead is done in COMICS [15]. This, as mentioned in the paper, is because the DCM website reports gross download numbers, not caring about single account downloads. This means, as reported in one comment on the website [^10] Comment on the most downloaded book Wanted Comics 11 -JVJ, that repeatedly downloads by automatic bots count as single downloads, thus invalidating the global score validity. Instead, we want to guarantee the presence of various style comics, featuring as many characters as possible and whose characters are mostly present in these selected books. Thus we propose an algorithm to select the books with the required principles: “Pow Selection Approach”, illustrated in the appendix of the main manuscript and reported also below.
D.1 Books selection algorithm
@@
16:         $book\_shared\_to\_unique\_ratio[book\_id]$leftarrow $ratio^{{\color[rgb]{0.4,0.6,0.8}pow}}$
@@
D.2 Comics books overview
@@
D.3 Offensive Content
Regarding offensive content, it is acknowledged that comics from the 1950s, both American and European, occasionally used terms that were mildly offensive towards minorities and wartime adversaries (such as Japanese or German soldiers), which were culturally tolerated during that period. However, following the implementation of the Comics Code in 1954 by the Comics Code Authority [^11] Comics Code History: The Seal of Approval https://cbldf.org/comics-code-history-the-seal-of-approval, content depicting offensive themes or inspiring violence was strictly prohibited. Consequently, our selection is confined exclusively to comics published post-1954, ensuring all included works are compliant with the Comics Code. This careful curation supports the use of the dataset in diverse research and educational settings without risking exposure to inappropriate material.
Appendix E Detailed Results
E.1 Detection
@@ E.2 Speaker identification
E.3 Character Naming and Dialog generation
@@
procedure EvaluateTranscription($model\_output,ground\_truth$)
@@
     $tot\_ed$, $char\_name\_score$ $\leftarrow 0,0,0$
@@
         $edit\_dist\leftarrow\text{calculate edit distance}(mo.text,gt.text)$
@@
         $anls\_score\leftarrow\text{calculate ANLS}(mo.name,gt.name)$
@@
     $char\_name\_score$leftarrow $char\_name\_score$ /// len($matches$)
@@
Table 9: Results on Character Naming. The metric reported is the ANLS on the names, which measures the similarity of two strings considering it wrong when more than 50% do not agree. The metric is calculated on every page and averaged over the CoMix datasets.
@@
Table 10: Results on Dialog generation. The metric reported is the minimum edit distance calculated for every possible match from ground truth and predicted dialogs. The metric is calculated on every page and averaged over the CoMix datasets
@@
Appendix F Copyrights information
Comics books, like many forms of art, are often subject to restricted access due to copyright laws. The images utilized for creating the CoMix dataset come from both existing datasets (eBDtheque, PopManga) and sources that are freely accessible (DCM, Comics).
In light of these considerations, we have developed a framework to systematically reorganize the image folders and filenames to align with our specified format. To utilize this dataset effectively, users should initially acquire the images from PopManga and eBDtheque directly from their respective sources. Following this, users can download our freely accessible images from DCM and Comics, along with our test annotations.
By employing the tools provided in our repository, users will be able to reformat the dataset according to our structure and replicate the results presented in our studies. This structured approach ensures compliance with copyright norms while facilitating ease of use and replicability of the dataset.

### v2

@@ 1 Introduction
In recent works[34], authors approached comics research to more complex tasks moving from classification[33], detection[36] and captioning[35] to diarization[19, 28], where low-level tasks - such as detecting objects, defining reading order and speaker identification and character re-identification - are used as a medium to generate an ordered transcription of who said what on single page. Despite the innovative application of this complex bottom-up approach to comics understanding tasks, there is no benchmark on character naming or diarization, nor metrics to assess their correctness [34].
@@
CoMix contains 3.8k images, gathered from 100 books, densely annotated with 130K objects across the four classes considered, 30k text-characters links, with 33k characters clusters with 16k total names identified. An overview of single-page annotations is provided in Figure 2 while statistics about datasets and annotations are provided in Table 2. We open-source the images and annotations of the validation splits. An evaluation server and images from the held-out test split are made available. Since currently there is no model that can tackle all the evaluation tasks in our benchmark, we provide baseline results for per-task models: object detection, speaker identification, character re-identification, reading order, character naming, and dialog generation.
@@ 3 Books in the CoMix
The CoMix dataset has been meticulously curated to showcase a broad spectrum of comic book styles, drawing samples from Japanese manga (PopManga), American comics (DCM and newly collected Comics), and French Bandes Dessinées (eBDtheque). Although multilingualism is not a primary objective, some pages from eBDtheque include French and Japanese, reflecting incidental multilingual aspects.
@@
Using this information, we selected comic books based on the distribution of character appearances across the books following the principle of (i) most characters possible, from the minimum number of books, and (ii) all possible instances of these characters should be in the 100 books selected, of which 20 goes to the test/val splits. In constructing CoMix, our primary goal is to underscore books featuring characters that recurrently appear across various publications, indicating significant narrative roles. To achieve this, the selection method computes the ratio of shared-to-unique character appearances for each book, thereby ranking and choosing the top books based on this ratio. Specifically, this ratio assesses the frequency of characters’ appearances within the same book against their appearances in other books. To further refine our selection, the algorithm was designed to maximize the diversity of characters across the selected books, prioritizing the retention of books that uniquely feature these characters. This ensures a broad representation of characters, with most being exclusive to the selected subset and not featured in the broader set of $22k$ comics. The ratio is then exponentiated (in our implementation, squared) to enhance distinctions between books with high and low levels of character sharing. After calculating these ratios for all books, they are ranked in descending order based on their scores, leading to the selection of the top 100 books. This metric, referred to as the “Pow Selection Approach”, is detailed in the Supplementary materials. This approach ensures that the selected books reflect broader narrative arcs and character interconnectivity, thereby enriching the CoMix dataset’s utility for diverse comic book analysis.
Splits: The CoMix contains 80 books from different sources, each of 44 pages on average, with a total of 3.5k images. This corresponds only to the held-out test split available through the evaluation server. Moreover, a validation split (20 books, 466 images), on which we tested and reported the results in this paper is provided.
@@ 4 Annotation in the CoMix benchmark
Table 2: Annotations in the CoMix benchmark. The reported numbers correspond to both test (80 books) and validation (20 books) splits. Each detected object is represented by the class and bounding box. Text elements also possess ground truth transcriptions. Character elements, whenever relevant, are associated with the name. Not all texts are associated with a character, but all the texts have Reading order ground truth.
@@
Dialog Generation. As a way to connect high-level and low-level single-page comics understanding capabilities, we define dialogs as an ordered list of tuples with the speaker name (character names or “narrator”). The texts are sorted based on the ground truth reading order. Annotators are instructed to maintain punctuations and text format (lower-case and upper-case) but discard other text properties (bold, italics, canceled, or underlined text).
@@ 5 Baselines results
Computational tasks: We defined six computational tasks based on the annotations available in the CoMix dataset. We summarise in Table 3 the task definitions (outputs, metrics) and the performance of top-performing baselines. It can be observed that the CoMix combines lower-level dense prediction tasks like object, speaker identification, and character re-identification, whose outputs are box and group of indexes, with higher-level tasks like dialog generation. More details about the task definitions are included in the supplementary materials.
Baselines: Ideally, a single model should be able to perform all the tasks in the CoMix benchmark. Since such a model is not available in the literature, we include results obtained with per-task baselines on the validation split for all the six tasks in the CoMix; see Table 3 for a summary of top-performing baselines and their average performance, and the supplementary materials for more details. When selecting and running these baselines, we favored the same approaches used by Sachdeva and Zisserman [28], which we carefully detail in the following section. However, for character naming, such models do not exist in the literature, so we evaluated namings together with dialogs, instead of with detection and clustering.
@@
Object detection: For object detection, we selected two common metrics: mean Average Precision at IoU of $0.5$ ($mAP@0.5$) and Recall at 100 objects ($R@100$). We benchmarked a variety of models including convolutional and transformer-based architectures, with a focus on their performance in both fine-tuned and zero-shot settings on comic data. Among these, GroundingDino [20], designed for open-set object detection using natural language object classes, was a key zero-shot model utilized to detect four classes: panels, characters, text, and faces. Additionally, conventional models such as Faster R-CNN, SSD, and YOLO were trained on varied comic styles to assess the impact of different training data distributions. For characters and faces detection we employed also a YOLOX-based available model named DASS [32]. Notably, Magi [28], a transformer-based model, was included for its impressive capability demonstrated in Manga-style comics. Details about the fine-tuned training procedure and GroundingDino prompts for zero-shot detection are provided in the appendix, as well as more detailed benchmark results.
@@
Character Naming & Dialog Generation. In addressing the challenges of character naming and dialog generation, we acknowledge that there is a lack of specific metrics. In previous works, the dialog is generated with a combination of algorithmic panel ordering, character progressive naming (“character 1”, “character 2”), and fine-tuned OCR for textbox transcription. This engineering and multi-step approach is a good showcase but is not evaluated with some metric against any ground truth. We introduce a smoothed case-sensitive edit distance metric called “Hybrid Dialog Score” ($HDS$) to evaluate the accuracy of model outputs against ground truth dialog annotations. This metric assesses both the precision of transcribed dialogues and the accuracy of character identification in a unified framework. The metric operates in three steps: (i) we match ground truth and predictions texts using Hungarian matching with edit distance metric; for every match, we calculate (ii) the textbox edit distance (normalized on ground truth text length) and (iii) the character name ANLS. An in-depth overview of the metric is provided in the supplementary materials, together with pseudocode. For our benchmarks, we utilized Magi as described in [28]. As MAGi is not able to detect names, we employ GPT-4 [26], leveraging its multimodal capabilities to interpret comic pages and generate structured outputs that include both dialogues and character names. GPT-4 was tasked with identifying and naming characters, where known, or assigning incremental identifiers otherwise, and transcribing dialogues exactly as they appear in comic format, respecting the sequence and case sensitivity. This approach not only captures the complexity of comic narratives but also enhances the evaluation of dialog transcription fidelity and character consistency across various comic styles. The result is reported in Table 3, and an overview of the GPT-4 prompt to obtain structured predictions is given in supplementary materials.
6 Ethical Considerations
Copyrights and Consent.
Comics, as a form of artistic expression, are governed by copyright laws that restrict access and usage. Our dataset, CoMix, aggregates comics from diverse sources: American comics from the Digital Comic Museum, manga from PopManga, and French Bande Dessinée from eBDtheque. In particular, Digital Comic Museum contains public domain assets, either released without copyright or with expired rights, allowing free research use [21]. The PopManga images are publicly accessible on “Manga Plus by Shueisha” with official permissions from copyright holders. The eBDtheque provides publicly available data cleared for non-commercial research use.
To ensure compliance, we have reorganized the dataset’s structure, enabling users to acquire images directly from original sources and utilize our repository tools for formatting and validation [36]. This approach maintains adherence to copyright norms while promoting dataset accessibility and replicability.
Data Quality and Representativeness.
The CoMix benchmark evaluates models across various comic styles—American, Japanese, and European—to represent major comic production hubs. Predominantly comprising out-of-copyright works, especially American comics from the 1950s, the dataset may inherently reflect the social biases and stereotypes of that era. To address potential biases, we ensure a balanced representation of comic styles and origins to minimize cultural bias (Diverse Dataset Composition) and incorporate statistical analysis of factors like gender, ethnicity, and language representation (Bias Detection and Analysis).
On this regard, our first analysis concerns biases related to appearance (color vs. black-and-white) and types of characters in the comics. The color statistics are straightforward: by examining the first pages, we could determine whether each comic chapter is in color or black-and-white, as shown in Table 4(a). For character type analysis, we classified character crops into one of four categories listed in Table 4(b), using a state-of-the-art open-source MLLM [^5] MiniCPM-llama3-v-2.5 at https://huggingface.co/openbmb/MiniCPM-Llama3-V-2_5), which is known for its strong visual recognition and instruction following capabilities.
While the dataset shows a bias towards male characters, this is considered a reflection of actual character type distributions in real comics and manga, rather than a flaw in the dataset’s quality. This dataset is intended for tasks such as dialog transcription testing.
Table 4: Initial Statistics in the CoMix Dataset
(a) Color vs. black-and-white images.
Type Percentage
Color 59.2%
Black-and-White 40.8%
(b) Character Types in the CoMix Dataset
Character Type Percentage
Male 74.3%
Female 17.9%
Animals 6.3%
Other 1.5%
Future iterations will expand to include underrepresented styles such as Webtoons and Manhwa, alongside additional languages, to enhance global applicability and reduce cultural over-representation. Detailed statistics and bias analyses are provided in the supplementary materials.
Semantic Harmful Content.
Our dataset has been meticulously curated to exclude NSFW and offensive content. All included works comply with the Comics Code Authority guidelines, ensuring appropriateness for diverse research and educational purposes. We conducted an automatic analysis of semantic content using the Llama3-80B model to classify panel-level text for offensiveness. Initial findings indicate low levels of harmful content (less than 0.5%), with detailed results and model prompts available in the supplementary materials.
Overall, these ethical considerations ensure that our research adheres to high standards of integrity, fairness, and respect for diverse cultural norms [21].
@@ 7 Conclusion
We introduce CoMix, a novel benchmark for multi-task and multi-modal comic analysis that addresses the limitations of existing datasets by incorporating diverse comic styles—including American, manga, and French—and providing comprehensive annotations across a wide range of tasks. CoMix encompasses fundamental vision tasks such as object detection and character re-identification, alongside complex multi-modal reasoning tasks like character naming and dialogue generation. The introduction of the Hybrid Dialog Score offers innovative metrics for evaluating these advanced tasks.
Baseline evaluations reveal significant performance gaps between state-of-the-art models and human performance, highlighting the inherent challenges in achieving nuanced understanding of the interplay between visual and textual elements in comics. By releasing the CoMix validation split and establishing an evaluation server for the held-out test split, we promote open research and facilitate robust benchmarking. CoMix sets a new standard for comprehensive comic analysis, providing a diverse and challenging testbed that will drive the development of more sophisticated and generalizable models capable of human-like comprehension in this culturally rich medium.
Acknowledgments and Disclosure of Funding
This paper has been supported by the Consolidated Research Group 2021 SGR 01559 from the Research and University Department of the Catalan Government, and by project PID2023-146426NB-100 funded by MCIU/AEI/10.13039/501100011033 and FSE+. This work has been also funded by the European Lighthouse on Safe and Secure AI (ELSA) from the European Union’s Horizon Europe program under grant agreement No 101070617.
Moreover, we are grateful to Digital Comics Museum for providing a free accessible source of comic books and to Grand Comics Database for the effort of building and maintaining correct and updated versions of comic memories. We want to thank Mohammed Ali Souibgui and Marco Mistretta for providing insightful input and Niccolò Biondi, Irene Campaioli, and Mariateresa Nardoni for being part of the revision and annotation team.
Supplementary Material
@@
This supplementary document complements the main paper by providing additional information and examples that could not be included within the page constraints of the original manuscript. The structure of this document is as follows: in Section S1 details about the models and their usage. In Section S2 we presents an overview of the annotations from existing datasets and compares them with those in the CoMix benchmark. In Section S3 we detail the selection process of American Comics books included in our study. Extensive results are provided in Section S4, including those reported in the main manuscript. Finally, copyright and biases information is outlined in the Ethical Section S5. This supplementary material is intended to enhance the understanding and transparency of the research presented in the main paper.
S1 Detailed Model Descriptions and Settings
@@
S2 Annotations overview
@@
Additionally, annotations for speaker identification, character re-identification, and reading order may be less visible in the images provided. These are represented as green and red lines connecting the text to the character boxes (speaker identification), colored points identifying each character bounding box (character re-identification), and purple splines connecting the text boxes in reading order orientation (reading order). For a clearer view of these annotations, we will release the validation set images and annotations on the project website [^8] Repository link: https://github.com/emanuelevivoli/CoMix-dataset..
@@
S3 Data selection
The CoMix benchmark sees the presence of selected American golden-age comics from DCM  [^9] Digital Comics Museum at https://digitalcomicmuseum.com. These documents have not been selected following the “most downloaded” principle, as instead is done in COMICS [15]. This, as mentioned in the paper, is because the DCM website reports gross download numbers, not caring about single account downloads. This means, as reported in one comment on the website [^10] Comment on the most downloaded book Wanted Comics 11 -JVJ, that repeatedly downloads by automatic bots count as single downloads, thus invalidating the global score validity. Instead, we want to guarantee the presence of various style comics, featuring as many characters as possible and whose characters are mostly present in these selected books. Thus we propose an algorithm to select the books with the required principles: “Pow Selection Approach”, reported in the Algorithm 1.
S3.1 Books selection algorithm
@@
16:         $book\_shared\_to\_unique\_ratio[book\_id]$leftarrow $ratio^{{\color[rgb]{0.4,0.6,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.6,0.8}pow}}$
@@
S3.2 Comics books overview
@@
S4 Detailed Results
S4.1 Detection
@@
S4.3 Character Naming and Dialog generation
@@
3:     $tot\_ed$, $char\_name\_score$ $\leftarrow 0,0,0$
@@
7:         $anls\_score\leftarrow\text{calculate ANLS}(mo.name,gt.name)$
@@
Table 10: Results on Character Naming and Dialog Generation. The metrics reported are ANLS for Character Naming and minimum edit distance for Dialog Generation. Both metrics are calculated on every page and averaged over the CoMix datasets.
Character Naming. The ANLS metric measures the similarity of two strings, considering it wrong when more than 50% do not agree.
@@
Dialog Generation. The minimum edit distance metric is calculated for every possible match between ground truth and predicted dialogs.
@@
S5 Additional Ethical Considerations
Alongside the copyright and consent information and the initial biases analysis featured in the main paper, we include information on automatic semantic harmful content analysis using cutting-edge Multimodal LLMs.
S5.1 Semantic Harmful Content
Regarding offensive content, it is acknowledged that comics from the 1950s, both American and European, occasionally used terms that were mildly offensive towards minorities and wartime adversaries (such as Japanese or German soldiers), which were culturally tolerated during that period. However, following the implementation of the Comics Code by the Comics Code Authority [^11] https://cbldf.org/comics-code-history-the-seal-of-approval, content depicting offensive themes or inspiring violence was strictly prohibited. Consequently, our selection is confined exclusively to comics published post 1954, ensuring all included works are compliant with the Comics Code. This careful curation supports the use of the dataset in diverse research and educational settings without risking exposure to inappropriate material.
Moreover, performing a full manual analysis of the semantic content and possible biases present in the CoMix dataset is complicated, and prone to subjective biases of human evaluators. We have instead opted to perform an automatic analysis of the textual content at the panel level, using the Llama3-80B model. We provided the model with detailed descriptions of different semantic classes and classified each panel accordingly (see the classes in Table 11).
The model was provided with detailed descriptions of different semantic classes, as shown below:
Model Prompt You are an image classifier trained to identify the offensiveness of comic balloon content. You will be provided with the extracted content of a panel, in a single row, which represents the concatenation of multiple balloons, following the reading order. Based on that, you must choose a number from -1 to 5 as one of the following: <TABLE 11> Examine the following text and determine the offensiveness score (-1, 0, 1, …, 5). Respond only with the identified style, without any explanation. <text>
Table 11: Classes provided to the model. The critical panels are the ones classified as 3, 4, and 5.
Code Class Short Description
-1 Empty The text is either empty or provides content which meaning is ‘empty text’.
0 Neutral Completely neutral content with no offensiveness.
1 Informal Casual content, maybe mild slang but not offensive.
2 Humorous/Sarcastic Humorous content, unlikely to offend most people.
3 Sexual/Angry/Violent Sexual, angry, or violent content, potentially inappropriate for some but not offensive.
4 Offensive Insulting content, but not racist or sexist. Just clearly offensive.
5 Highly Offensive/Racist/Sexist Extremely racist, sexist, or offensive.
From the initial analysis, we were able to obtain the results shown in Table 12.
Table 12: Semantic harmful content analysis.
code class percentage %
-1 Empty 23.9 %
0 Neutral 13.3 %
1 Informal 14.5 %
2 Humorous/Sarcastic 34.2 (34.7) %
3 Sexual/Angry/Violent 12.7 (13.1) %
4 Offensive 1.6 ( 1.2 ) %
5 Highly Offensive/Racist/Sexist 0.8 (0.3) %
We subsequently manually checked 370 panels to verify these results (from classes 3,4,5). We found that many of these are actually less (or not) offensive than the model classifications. An example of a panel being classified as “Highly Offensive/Racist/Sexist” is provided in Figure 10.
Figure 10: Example of a “Highly Offensive/Racist/Sexist” image (based on textual information) that, instead, belongs to “violent” or “incitement to violence” class.
We updated the percentage in Table 12 considering these manual verifications. These statistics are meant to be indicative of the semantic content of our dataset, while we consider that a full, detailed study is out of the context of this manuscript.